# Controlled Retrieval-augmented Context Evaluation for Long-form RAG

### v1

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models by incorporating context retrieved from external knowledge sources. While the effectiveness of the retrieval module is typically evaluated with relevance-based ranking metrics, such metrics may be insufficient to reflect the retrieval’s impact on the final RAG result, especially in long-form generation scenarios. We argue that providing a comprehensive retrieval-augmented context is important for long-form RAG tasks like report generation and propose metrics for assessing the context independent of generation. We introduce CRUX, a Controlled Retrieval-aUgmented conteXt evaluation framework designed to directly assess retrieval-augmented contexts. This framework uses human-written summaries to control the information scope of knowledge, enabling us to measure how well the context covers information essential for long-form generation. CRUX uses question-based evaluation to assess RAG’s retrieval in a fine-grained manner. Empirical results show that CRUX offers more reflective and diagnostic evaluation. Our findings also reveal substantial room for improvement in current retrieval methods, pointing to promising directions for advancing RAG’s retrieval. Our data and code are publicly available to support and advance future research on retrieval. [^1] https://anonymous.4open.science/r/rag-rerank-85CF
@@
Jia-Huei Ju1  Suzan Verberne2 Maarten de Rijke1  Andrew Yates3 1University of Amsterdam 2Leiden University 3Johns Hopkins University
@@ 1 Introduction
With their emerging instruction-following capabilities Ouyang et al. (2022); Wei et al. (2021), large language models (LLMs) have adopted retrieval-augmented generation (RAG) Lewis et al. (2020); Guu et al. (2020) to tackle more challenging tasks, such as ambiguous question answering (QA) Stelmakh et al. (2022); Gao et al. (2023) or long-form response generation Shao et al. (2024). The role of retrieval in RAG is to access information from external sources and prompt it as plug-in knowledge for LLMs. To achieve this, typical RAG systems retrieve the $k$ most relevant chunks as the retrieval-augmented context (abbreviated as retrieval context, hereafter), and prompt the LLM to generate a response using this information.
@@
Figure 1: An example of long-form generation on CRUX with open-ended query $x$ and desired response $y$. The underlined text marks relevant content in the retrieval (Refer to caption) that contributes to the final result. By directly assessing retrieval context $Z$, we can further explicitly identify incomplete (Refer to caption) and redundant retrieval (Refer to caption).
@@
It was found that a suboptimal retrieval context hinders the generation process Asai et al. (2024); Rau et al. (2024), triggering negative impacts and resulting in unsatisfying final RAG results. One of widely-studied effects is the impact of noise from irrelevant retrieval Yoran et al. (2023), which increases the risk of hallucinations Asai et al. (2022) and distractions Shi et al. (2023). Such prior studies have mainly focused on short-answer tasks; however, recent RAG development has shifted towards generating comprehensive and structured reports with open-ended queries Zhao et al. (2024); Lawrie et al. (2024), as illustrated in Figure 1, introducing new concerns of suboptimal retrieval.
@@
In the scenario of open-ended queries where a short answer is insufficient and a long-form result is required, incompleteness and redundancy emerge as the critical yet underexplored negative impacts from retrieval Joren et al. (2024). Specifically, (i) incomplete retrieval fails to capture the full nuance of the query, leading to partial or misleading generations. (ii) Redundant retrieval contexts restrict the diversity of knowledge, undermining the usefulness of augmented knowledge Yu et al. (2024); Chen and Choi (2024). Figure 1 exemplifies such impacts of suboptimal retrieval matters on the final long-form RAG result.
@@
To examine these effects, a suitable retrieval evaluation framework is crucial for measuring completeness and redundancy in the retrieval context. Current retrieval evaluation approaches are insufficient for measuring retrieval effectiveness in long-form RAG, as they are designed for web search Bajaj et al. (2016) or short-answer QA Kwiatkowski et al. (2019). They only require a focus on relevance-based ranking, which can be simply evaluated with retrieval metrics such as MRR and Recall@$k$. In contrast, long-form RAG requires retrieving multiple aspects and subtopics to ensure completeness, which goes beyond surface-level relevance Tan et al. (2024); Grusky et al. (2018).
@@
To address the gap, we propose a Controllable Retrieval-aUgmented conteXt evaluation framework (CRUX). The framework includes controlled evaluation datasets and coverage-based metrics, which directly assess the content of the retrieval context instead of relevance-based ranking. We use human-written multi-document summaries to define the scope of retrieval context, enabling a controlled oracle retrieval for more diagnostic evaluation results. Finally, we assess both (intermediate) retrieval context and (final) RAG result via question-based evaluation Sander and Dietz (2021), supporting fine-grained and more aligned evaluation between them.
@@
To validate the usability of our evaluation framework, we conduct empirical experiments with multiple retrieval and re-ranking strategies, including relevance and diversity re-ranking. Empirical results explicitly reveal the limitations of suboptimal retrieval in terms of coverage and density. Our additional metric analysis further demonstrates that relevance ranking metrics lack coverage-awareness, highlighting CRUX’s strength in identifying retrieval impacts on long-form RAG. Notably, our framework balances scalability and reliability by integrating LLM-based judgments with human-grounded data. Our final human evaluation also confirms CRUX’s alignment with human perception.
@@
  • We create a controlled dataset tailored for evaluating retrieval context for long-form RAG;
  • We propose coverage-based metrics with upper bounds to help diagnosing retrieval context in terms of completeness and redundancy;
@@ 2 Related Work
Evaluation methodologies in IR and NLP have been standardized and developed for decades Voorhees (2002, 2004). In recent years, nugget-based (sub-topics or sub-questions) evaluation Pavlu et al. (2012); Clarke et al. (2008); Dang et al. (2008) has resurfaced as an important focus due to the feasibility of automatic judgments. Similarly, question-based evaluation estimate the answerability Eyal et al. (2019); Sander and Dietz (2021) of given text, is well-aligned with LLMs while preserving aspect-level granularity, making it particularly good for evaluating long-form generation. This helps to inform the development of a unified evaluation setup for both intermediate retrieval context and final long-form results, thereby facilitating more informative evaluation for RAG’s retrieval methods.
@@ 3 Controlled Retrieval-augmented Context Evaluation (CRUX)
This section introduces CRUX, a controlled evaluation framework for assessing retrieval context in long-form RAG. It comprises: (1) definitions of retrieval context and its sub-question answerability (§ 3.1); (2) curated evaluation datasets (§ 3.2) and (3) answerability-driven performance metrics: coverage and density (§ 3.3).
@@ 3.1 Retrieval-augmented Context
$y\leftarrow G(x,Z,I),\quad Z\leftarrow RA_{\theta}(x,\mathcal{K}).\$ (1)
$RA_{\theta}$ denotes the retrieval modules that augment retrieval context $Z$ from an external knowledge source $\mathcal{K}$ (i.e., a corpus), and $G$ is a LLM generator that input with the query $x$, retrieval context $Z$ and a task-specific instruction prompt $I$, to generate the final long-form RAG result $y$. Particularly, we argue that the quality of retrieval context is a key limitation for achieving optimal RAG results and propose an evaluation framework to diagnose it.
@@
To assess retrieval context quality beyond relevance-based ranking, we adopt question-based evaluation Eyal et al. (2019); Sander and Dietz (2021). We assess the content of an arbitrary text $z$ with a diverse set of knowledge-intensive sub-questions $Q=\{q_{1},q_{2},\ldots,q_{n}\}$. Such diversity enables these questions to serve as a surrogate for evaluating multiple aspects of a query, thereby facilitating explicit diagnosis of underlying concerns such as completeness and redundancy. Specifically, we use an LLM to judge how well the text $z$ answers each sub-question and estimate a binary sub-question answerability value (answerability, hereafter):
@@ 3.2 Data Creation for Controlled Evaluation
We further construct datasets tailored for our evaluation framework to support controlled analysis. As illustrated in Figure 2, we treat human-written multi-document summaries as the central anchor for defining: (1) the explicit scope of relevant retrieval context $Z^{*}$; (2) an open-ended query $x$; (3) a diverse set of sub-questions $Q$. Together, these components support our assessment of completeness and redundancy.
Explicit scope of retrieval context.
@@
We use an LLM to synthesize a query with open-ended information needs from the human-written summary $y^{*}$ via in-context prompting Brown et al. (2020) (See an example in Figure 1 and 9 also). We denote these queries as $x$ in Eq. (1), which is the initial input for both retrieval and generation. Such queries help expose limitations in existing retrieval systems, which often return either irrelevant or redundant passages, resulting in incomplete retrieval contexts. Notably, the query generation process is adaptable and can be tailored to various kinds of queries Yang et al. (2024) via similar in-context prompting.
@@
Similarly, we synthesize a diverse set of knowledge-intensive sub-questions $Q$ from the human-written summary which cover the highlights in the oracle RAG results (i.e., $y^{*}$). Thanks to the controlled settings, for each query $x$, we enumerate all possible pairs of sub-questions $q\in Q$ and relevant passages $p\in P^{*}$, then judge them with an LLM. Hence, for each relevant passage, we obtain a list of graded ratings for all the sub-question as mentioned in Eq. (2). Finally, we can obtain the matrix of graded ratings as shown in Figure 2. In addition, the judged ratings can serve as consistency filtering to identify unanswerable sub-questions for mitigating out-of-scope and hallucinated questions. These pre-judged ratings can be further reused for evaluating retrieval context, which is also released with the data.
@@
Once we have pre-judgements of all relevant passages $p\in P^{*}$, we further identify which passages are truly necessary and construct a smaller subset of relevant passages, denoted as $P^{**}$. Specifically, we define this required subset if the passages in the subset can collectively answer all sub-questions $q\in Q$. To do so, we first rank each relevant passage according to how many questions it can answer and greedily assigned it to the subset until no additional sub-questions can be answered. [^3] The default answerability threshold $\eta$ is set to 3. The remaining are categorized as either partially or fully redundant.
@@
Due to the limited computational resources, we finally collected 100 open-ended queries from Multi-News Fabbri et al. (2019) and 50 queries from DUC Over and Yen (2004). The knowledge source $\mathcal{K}$ has around 500K passages, collected from training and test splits of Multi-News and the DUC. We generate all data using open-source Llama-3.1-70B-Instruct. [^4] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct Meta (2024). Detailed data statistics and generation settings are reported in Appendix A.1.
@@ 3.3 Evaluation Metrics
We define three performance metrics to assess retrieval context for long-form RAG. We begin by measuring context’s completeness using coverage, then introduce derived metrics: ranked coverage and density to further take redundancy into account.
Figure 3: CRUX employs sub-question answerability to directly assess the textual content of both retrieval context $Z$ and its corresponding RAG result $y$. The metrics include coverage and density.
@@
Rather than evaluating the retrieval results based on only their relevance (e.g., nDCG and MAP), we assess the content of the retrieval contexts based on answerability. Given a retrieval context $Z$, we explicitly quantify the context’s coverage with how many questions it can answer over the answerable sub-questions. To compute this, we aggregate graded ratings by taking the maximum across passages in retrieval context $Z$ and obtain binary answerability as depicted in Figure 3. We finally normalize it by the total number of answerable sub-questions. Formally, the coverage of the retrieval context is defined as:
@@
[…]
$ng(r)=$ $\sum^{|Q|}_{i=1}\mathcal{I}_{i,r}(1-\alpha)^{c_{i,r-1}},$ (4)
where $r$ is the passage rank position in the retrieval context. The function $ng$ is novelty gain, representing how much new information is covered with respect to the position $r$ and sub-question $q_{i}$. Discount factor $\alpha$ is used for penalizing redundant sub-questions when accumulating gains.
@@
Table 1: Evaluation results of empirical retrieval contexts $Z$ and corresponding final results $y$ (the columns in gray) on CRUX-DUC and Multi-News. Scores with bold font and underlined are the highest and lowest. For each dataset, columns 1 and 2 show retrieval coverage and ranked coverage; column 3 shows the final result coverage. The last two columns are density of retrieval context and final result. The bottom row reports the ranking correlation between retrieval context and final results.
@@ 4.1 Experimental Setups
Our experiments employ varying cascaded retrieval pipelines to augment context from the knowledge corpus. Given an open-ended query $x$, we first retrieve top-100 relevant candidate passages. Three initial retrieval approaches are considered: lexical retrieval (LR) with BM25, [^5] https://github.com/castorini/pyserini/ dense retrieval (DR) and learned sparse retrieval (LSR) using ContrieverFT Izacard et al. (2021) and SPLADE-v3 Lassance et al. (2024).
@@
We further re-rank the 100 candidate passages with more effective models, constructing the final retrieval context $Z$. We experiment with varying re-ranking strategies, including pointwise re-ranking models (#parameters): miniLM (220M) and monoT5 (3B). In addition, we include state-of-the-art LLM-based listwise re-ranking models: RankZephyr (7B) Pradeep et al. (2023) and RankFirst (7B) Reddy et al. (2024), as well as the Setwise re-ranking (3B) Zhuang et al. (2023). Lastly, we evaluate the maximal marginal relevance (MMR) algorithm for diversity re-ranking to consider both relevance and diversity. [^6] We follow Gao and Zhang (2024) and adopt the same pre-trained encoder for MMR: https://huggingface.co/sentence-transformers/all-mpnet-base-v2
@@
As our goal is to analyze how incomplete and redundant retrieval context affects the final RAG result, we assess both the quality of retrieval context $Z$ and further investigate the relationships between them and final coverage and density: $Cov(y)$ and $Den(y)$. Notably, the explicit scope of relevant passages allows us to reuse the pre-judgements for relevant passages as shown in Figure 3. Unless otherwise specified, we set the default answerability threshold $\eta$ to 3.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models by incorporating context retrieved from external knowledge sources. While the effectiveness of the retrieval module is typically evaluated with relevance-based ranking metrics, such metrics may be insufficient to reflect the retrieval’s impact on the final RAG result, especially in long-form generation scenarios. We argue that providing a comprehensive retrieval-augmented context is important for long-form RAG tasks like report generation and propose metrics for assessing the context independent of generation. We introduce CRUX, a Controlled Retrieval-aUgmented conteXt evaluation framework designed to directly assess retrieval-augmented contexts. This framework uses human-written summaries to control the information scope of knowledge, enabling us to measure how well the context covers information essential for long-form generation. CRUX uses question-based evaluation to assess RAG’s retrieval in a fine-grained manner. Empirical results show that CRUX offers more reflective and diagnostic evaluation. Our findings also reveal substantial room for improvement in current retrieval methods, pointing to promising directions for advancing RAG’s retrieval. Our data and code are publicly available to support and advance future research on retrieval for RAG. [^1] https://github.com/DylanJoo/crux
@@
Jia-Huei Ju1  Suzan Verberne2  Maarten de Rijke1  Andrew Yates3 1University of Amsterdam  2Leiden University 3Johns Hopkins University, HLTCOE {j.ju, m.derijke}@uva.nl,  s.verberne@liacs.leidenuniv.nl, andrew.yates@jhu.edu
@@ 1 Introduction
With their emerging instruction-following capabilities Ouyang et al. (2022); Wei et al. (2021), large language models (LLMs) have adopted retrieval-augmented generation (RAG) Lewis et al. (2020); Guu et al. (2020) to tackle more challenging tasks, such as ambiguous question answering (QA) Stelmakh et al. (2022); Gao et al. (2023) and long-form response generation Shao et al. (2024). The role of retrieval in RAG is to access information from external sources and prompt it as plug-in knowledge for LLMs. To achieve this, typical RAG systems retrieve the $k$ most relevant chunks as the retrieval-augmented context (abbreviated as retrieval context, hereafter), and prompt the LLM to generate a response using this information.
@@
Figure 1: An example of long-form generation with an open-ended query $x$ and a desired response $y$. The underlined text marks relevant content in the retrieval (Refer to caption) that contributes to the final result. By directly assessing the retrieval context $Z$, we can further explicitly identify incomplete (Refer to caption) and redundant retrieval (Refer to caption).
@@
A suboptimal retrieval context hinders the generation process Asai et al. (2024); Rau et al. (2024), triggering negative impacts and resulting in unsatisfying final RAG results. One widely-studied effect is the impact of noise from irrelevant retrieval Yoran et al. (2023), which increases the risk of hallucinations Asai et al. (2022) and distractions Shi et al. (2023). Such prior studies have mainly focused on short-answer tasks; however, recent RAG research has shifted towards generating comprehensive and structured reports with open-ended queries Zhao et al. (2024); Lawrie et al. (2024), as illustrated in Figure 1, introducing new concerns of suboptimal retrieval.
@@
In the scenario of open-ended queries where a short answer is insufficient and a long-form result is required, incompleteness and redundancy emerge as the critical yet underexplored negative impacts from retrieval Joren et al. (2024). Specifically, (i) incomplete retrieval fails to capture the full nuance of the query, leading to partial or misleading generations, and (ii) redundant retrieval contexts restrict the diversity of knowledge, undermining the usefulness of augmented knowledge Yu et al. (2024); Chen and Choi (2024). Figure 1 exemplifies such impacts of suboptimal retrieval matters on the final long-form RAG result.
@@
To examine these effects, a suitable retrieval evaluation framework is crucial for measuring completeness and redundancy in the retrieval context. Current retrieval evaluation practices are insufficient for measuring retrieval effectiveness in long-form RAG, as they are designed for web search Bajaj et al. (2016) or short-answer QA Kwiatkowski et al. (2019). They only require a focus on relevance-based ranking, which can be simply evaluated with retrieval metrics such as MRR and Recall@$k$. In contrast, long-form RAG requires retrieving multiple aspects and subtopics to ensure completeness, which goes beyond surface-level relevance Tan et al. (2024); Grusky et al. (2018).
@@
To address the gap, we propose a Controllable Retrieval-aUgmented conteXt evaluation framework (CRUX). The framework includes controlled evaluation datasets and uses coverage-based metrics that directly assess the content of the retrieval context instead of relevance-based ranking. We use human-written multi-document summaries to define the scope of the retrieval context, enabling a controlled oracle retrieval for more diagnostic evaluation results. Finally, we assess both the (intermediate) retrieval context and (final) RAG result via question-based evaluation Sander and Dietz (2021); Dietz (2024), supporting fine-grained evaluation between them.
@@
To validate the usability of our evaluation framework, we conduct empirical experiments with multiple retrieval and re-ranking strategies, including relevance and diversity re-ranking. Empirical results demonstrate the limitations of suboptimal retrieval in terms of coverage and density. Our additional metric analysis further demonstrates that relevance ranking metrics lack coverage-awareness, highlighting CRUX’s strength in identifying retrieval impacts on long-form RAG. Notably, our framework balances scalability and reliability by integrating LLM-based judgments with human-grounded data. Our final human evaluation also confirms CRUX’s alignment with human perception.
@@
  • We create a controlled dataset tailored for evaluating the retrieval context for long-form RAG;
  • We propose coverage-based metrics with upper bounds to help diagnosing the retrieval context in terms of completeness and redundancy;
@@ 2 Related Work
Evaluation methodologies in IR and NLP have been standardized and developed for decades Voorhees (2002, 2004). In recent years, nugget-based (sub-topics or sub-questions) evaluation Pavlu et al. (2012); Clarke et al. (2008); Dang et al. (2008) has resurfaced as an important focus due to the feasibility of automatic judgments. Similarly, question-based evaluation that estimates the answerability Eyal et al. (2019); Sander and Dietz (2021) of a given text is well-aligned with LLMs while preserving aspect-level granularity, making it particularly good for evaluating long-form generation. This helps inform the development of a unified evaluation setup for both the intermediate retrieval context and final long-form results, thereby facilitating more informative evaluation for RAG’s retrieval methods.
@@ 3 Controlled Retrieval-augmented Context Evaluation (CRUX)
This section introduces CRUX, a controlled evaluation framework for assessing the retrieval context in long-form RAG. It comprises: (1) definitions of retrieval context and its sub-question answerability (§ 3.1); (2) curated evaluation datasets (§ 3.2) and (3) answerability-driven performance metrics: coverage and density (§ 3.3).
@@ 3.1 Retrieval-augmented Context
$y\leftarrow G(x,Z,I),\quad Z\leftarrow RA_{\theta}(x,\mathcal{K}).\\$ (1)
$RA_{\theta}$ denotes the retrieval modules that augment the retrieval context $Z$ from an external knowledge source $\mathcal{K}$ (i.e., a corpus), and $G$ is a LLM generator that takes as input a query $x$, retrieval context $Z$, and task-specific instruction prompt $I$ to generate the final long-form RAG result $y$. Particularly, we argue that the quality of the retrieval context is a key limitation for achieving optimal RAG results and propose an evaluation framework for it.
@@
To assess the retrieval context quality beyond relevance-based ranking, we adopt question-based evaluation Eyal et al. (2019); Sander and Dietz (2021). We assess the content of an arbitrary text $z$ with a diverse set of knowledge-intensive sub-questions $Q=\{q_{1},q_{2},\ldots,q_{n}\}$. Such diversity enables these questions to serve as a surrogate for evaluating multiple aspects of a query, thereby facilitating explicit diagnosis of underlying concerns such as completeness and redundancy. Specifically, we use an LLM to judge how well the text $z$ answers each sub-question and estimate a binary sub-question answerability value (answerability, hereafter):
@@ 3.2 Data Creation for Controlled Evaluation
We further construct datasets tailored for our evaluation framework to support controlled analysis. As illustrated in Figure 2, we treat human-written multi-document summaries as the central anchor for defining: (1) the explicit scope of the relevant retrieval context $Z^{*}$; (2) an open-ended query $x$; (3) a diverse set of sub-questions $Q$. Together, these components support our assessment of completeness and redundancy.
Explicit scope of the retrieval context.
@@
We use an LLM to synthesize a query with open-ended information needs from the human-written summary $y^{*}$ via in-context prompting Brown et al. (2020). Examples are shown in Figures 1 and 9. We denote these queries as $x$ in Eq. (1), which is the initial input for both retrieval and generation. Such queries help expose limitations in existing retrieval systems, which often return either irrelevant or redundant passages, resulting in incomplete retrieval contexts. Notably, the query generation process is adaptable and can be tailored to various kinds of queries Yang et al. (2024) via similar in-context prompting.
@@
Similarly, we synthesize a diverse set of knowledge-intensive sub-questions $Q$ from the human-written summary which cover the highlights in the oracle RAG results (i.e., $y^{*}$). Thanks to the controlled settings, for each query $x$, we enumerate all possible pairs of sub-questions $q\in Q$ and relevant passages $p\in P^{*}$, then judge them with an LLM. Hence, for each relevant passage, we obtain a list of graded ratings for all the sub-question as mentioned in Eq. (2). Finally, we can obtain the matrix of graded ratings as shown in Figure 2. In addition, the judged ratings can serve as consistency filtering to identify unanswerable sub-questions for mitigating out-of-scope and hallucinated questions. These pre-judged ratings can be further reused for evaluating the retrieval context, which is also released with the data.
@@
Once we have pre-judgments of all relevant passages $p\in P^{*}$, we further identify which passages are necessary and construct a smaller subset of relevant passages, denoted as $P^{**}$. Specifically, we define this required subset as those passages that can collectively answer all sub-questions $q\in Q$. To do so, we first rank each relevant passage according to how many questions it can answer and greedily assign each to the subset until no additional sub-questions can be answered. [^3] The default answerability threshold $\eta$ is set to 3. Those remaining are categorized as either partially or fully redundant.
@@
We collected 100 open-ended queries from Multi-News Fabbri et al. (2019) and 50 queries from DUC Over and Yen (2004). The knowledge source $\mathcal{K}$ has around 500K passages, collected from training and test splits of Multi-News and the DUC. We generate all data using open-source Llama-3.1-70B-Instruct. [^4] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct Meta (2024). Detailed data statistics and generation settings are reported in Appendix A.1.
@@ 3.3 Evaluation Metrics
We use three metrics to assess the retrieval context for long-form RAG. We begin by measuring a retrieval context’s completeness using coverage, then introduce derived metrics: ranked coverage and density to further take redundancy into account.
Figure 3: CRUX employs sub-question answerability to directly assess the textual content of both the retrieval context $Z$ and its corresponding RAG result $y$. The metrics include coverage and density.
@@
Rather than evaluating the retrieval results based on only their relevance (e.g., nDCG and MAP), we assess the content of the retrieval contexts based on answerability. Given a retrieval context $Z$, we explicitly quantify the context’s coverage with how many questions it can answer over the answerable sub-questions. To compute this, we aggregate graded ratings by taking the maximum across passages in the retrieval context $Z$ and obtain binary answerability as depicted in Figure 3. We finally normalize it by the total number of answerable sub-questions. Formally, the coverage of the retrieval context is defined as:
$Cov(Z)=\dfrac{\#\{q\in Q|\max\Big(G(p\in Z,q,I_{g})\Big)\geq\eta\}}{|Q|}.$
@@
$\mbox{}\hskip-5.69054pt\alpha\text{-nDCG}=$ $\sum_{r=1}^{|Z|}\frac{ng(r)}{\log(r+1)}/{\sum_{r=1}^{|Z^{*}|}\frac{ng^{*}(r)}{\log(r+1)}}$ (4)
$ng(r)=$ $\sum^{|Q|}_{i=1}\mathcal{I}_{i,r}(1-\alpha)^{c_{i,r-1}}$ (5)
where $r$ is the passage rank position in the retrieval context. The function $ng$ is novelty gain, representing how much new information is covered with respect to the position $r$ and sub-questions $q_{i}$. Discount factor $\alpha$ is used for penalizing redundant sub-questions when accumulating gains.
@@
Table 1: Evaluation results of empirical retrieval contexts $Z$ and corresponding final results $y$ (the columns in gray) on CRUX-DUC and Multi-News. Scores with bold font and underlined are the highest and lowest. For each dataset, columns 1 and 2 show retrieval coverage and ranked coverage; column 3 shows the final result coverage. The last two columns are the density of the retrieval context and final result. The bottom row reports the ranking correlation between the retrieval context and final results.
@@ 4.1 Experimental Setups
Our experiments employ varying cascaded retrieval pipelines to augment context from the knowledge corpus. Given an open-ended query $x$, we first retrieve the top-100 relevant candidate passages. Three initial retrieval approaches are considered: lexical retrieval (LR) with BM25, [^5] https://github.com/castorini/pyserini/ dense retrieval (DR) using ContrieverFT Izacard et al. (2021) and learned sparse retrieval (LSR) using SPLADE-v3 Lassance et al. (2024).
@@
We further re-rank the 100 candidate passages with more effective models, constructing the final retrieval context $Z$. We experiment with varying re-ranking strategies, including pointwise re-ranking models: miniLM (220M) and monoT5 (3B). In addition, we include state-of-the-art LLM-based listwise re-ranking models: RankZephyr (7B) Pradeep et al. (2023) and RankFirst (7B) Reddy et al. (2024), as well as Setwise re-ranking (3B) Zhuang et al. (2023). Lastly, we evaluate the maximal marginal relevance (MMR) algorithm for diversity re-ranking to consider both relevance and diversity. [^6] We follow Gao and Zhang (2024) and adopt the same pre-trained encoder for MMR: https://huggingface.co/sentence-transformers/all-mpnet-base-v2
@@
As our goal is to analyze how incomplete and redundant retrieval context affects the final RAG result, we assess both the quality of retrieval context $Z$ and further investigate the relationships between them and final coverage and density: $Cov(y)$ and $Den(y)$. Notably, the explicit scope of relevant passages allows us to reuse the pre-judgments for relevant passages as shown in Figure 3. Unless otherwise specified, we set the default answerability threshold $\eta$ to 3.